OpenStudy (anonymous):

Can someone help with 7th grade algebra? i understand, but i still need help

OpenStudy (anonymous):

sure

OpenStudy (anonymous):

thanks. variables and percentages are difficult fot me

OpenStudy (anonymous):

have u got any question with u ?

OpenStudy (anonymous):

no, im sorry

OpenStudy (anonymous):

Then how am I supposed to exlain...if do not have any questions

OpenStudy (anonymous):

i do! can't you just make up a problem? gooses >:(

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

x + 5 = 6 ... something like this?

OpenStudy (anonymous):

yeah. that would be 1. right?

OpenStudy (anonymous):

ok... we need to solve for x...for this we have to subtract anything which is on left side...in this case we have 5 ... so x +5 - 5 = 6 -5 ... u got it?

OpenStudy (anonymous):

i think so?

OpenStudy (anonymous):

:(

OpenStudy (anonymous):

ok, so i just thought that you went like "Ok, so it is x + 5 = 6, so i can flip it around, to where it makes 6 - 5 = x, and the answer would be one" thats just how i think :/

OpenStudy (anonymous):

ok so are you doing Pre-algebra? is tht wat ur school calls it?

OpenStudy (anonymous):

yes.... i'm trying to get to just plain algebra too because i am really slow when it comes to math

OpenStudy (anonymous):

ya same here but i rocked it when i took pre agebra all A's All year. idk why... anyways so u say u dont have a problem but im in a much bigger math class than that right now so i cant make up a problem.... so u give me one and i can help!

OpenStudy (anonymous):

um ok.... say if it was like 4t - 6 = 27 ¿:ı somethin like that. i will brb i need to go make some food... it will take like one minute

OpenStudy (anonymous):

ok imma rite the steps the solve that problem it EASY

OpenStudy (anonymous):

k back

OpenStudy (anonymous):

so the problem is 4t - 6 = 27 and ur trying to find the value of T. The first step is: 4t - 6 = 27 +6 +6 4t = 27 + 6 ^ 33 4t = 33 4 4 The fours cancel out so that t = 33 / 4 ^ t = 8.25 Get it kinda?

OpenStudy (anonymous):

uh...

OpenStudy (anonymous):

?

OpenStudy (anonymous):

whats thr +6 +6 thinggy?

OpenStudy (anonymous):

thats just a step u need to do so that you can find what t equals. Bc 4T - 6 = 27 so that to find t u have to find oh well nvm

OpenStudy (anonymous):

i'm sorry

OpenStudy (anonymous):

i'm stupid :/ i suck at ,ath

OpenStudy (awhite):

basically if you have b+2=3 and you want to solve to b, to get rid of the 2 on the right hand side of the ='s sign, you have to subtract 2 from both sides of the equation. so you have: b+2-2 = 3-2 b+0 = 3-2 = 1 so b=1

OpenStudy (anonymous):

that makes some sense

OpenStudy (awhite):

basically whatever you do to one side of the equation, you have to also do to the other side

OpenStudy (awhite):

so if you have 2*a = 10 and you need to solve for a then you need to divide both sides by 2 to get rid of the 2 multiplying the a, and you end up with 2*a/2 = 10/2 = 5 so a = 5

OpenStudy (anonymous):

¿:ı i understnad most of that

OpenStudy (awhite):

so could you solve for c in: 2*v +3 = 23 ?

OpenStudy (awhite):

sorry solve for v!

OpenStudy (anonymous):

um yes.. so .... uhh

OpenStudy (awhite):

where do you think you'd start?

OpenStudy (anonymous):

at the 2*v right? can you turn it around?

OpenStudy (awhite):

It would be easier to start by getting rid of the 3 on the right hand side.

OpenStudy (anonymous):

ok. so, because it is a positive you would subtract right?

OpenStudy (awhite):

yep

OpenStudy (anonymous):

so, 2*v +3- 3 = 23 ?

OpenStudy (awhite):

subtract from both sides :)

OpenStudy (anonymous):

oh yeah :) 2*v +3- 3 = 23- 3 would = 20 so far right?

OpenStudy (awhite):

yep

OpenStudy (anonymous):

k so what would you do next?

OpenStudy (awhite):

well then as you have 2*v = 20, you want to get rid of the 2

OpenStudy (anonymous):

um would you do the same thing? but ddividing?

OpenStudy (awhite):

exactly. divide through on both sides.

OpenStudy (anonymous):

k so... i'm sorry i suck at division. whatt would you divide by?

OpenStudy (awhite):

so because you want to get rid of the 2, you divide through by 2. ie: 2*v/2 = 20/2

OpenStudy (awhite):

does that make sense?

OpenStudy (anonymous):

yes.. but why would you keep the *v ?

OpenStudy (anonymous):

i'm sorry :/

OpenStudy (awhite):

you are solving for v, so you want to end up with v on one side, and everything else on the other side of the equation.

OpenStudy (anonymous):

oh. so... you would end up with 1 right? but on the right side it would be 10.

OpenStudy (awhite):

yeah, because v*1 = v :) so v = 10 is the answer

OpenStudy (anonymous):

sweet! thanks (: